40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 12 19:26:08 2012, Rijuparna Chakraborty, Update, , cavitymode scan 7x
    Reply  Thu Sep 13 07:36:41 2012, Steve, Update, General, restore conditions 
Message ID: 7378     Entry time: Thu Sep 13 07:36:41 2012     In reply to: 7376
Author: Steve 
Type: Update 
Category: General 
Subject: restore conditions 

Quote:

 Summary: Recorded the presence of higher order modes in IMC

What I did: Misaligned the flat mirror MC1 by small amount in both pitch and yaw (it was needed to be done cause at the beginning of the experiment no higher order modes were present)  and scanned the cavity for frequency-range 32MHz to 45MHz.

I found the presence of higher order modes around 36.7MHz (1st order)  and 40.6MHz (2nd order) along with two other strong modes near 35MHz and 42.5MHz.

 

 Please, restore condition after you finished and update elog right away! People wasted hours yesterday not knowing the condition of the MC

ELOG V3.1.3-